OpenStudy (anonymous):

Logarithmic equation? log3(x-4)=-3 so what does x=

OpenStudy (anonymous):

its x-4 = 1/27

OpenStudy (anonymous):

start by rewriting in equivalent exponential form

OpenStudy (anonymous):

now 27x- 108= 0

OpenStudy (anonymous):

\[\log_b(\heartsuit)=\clubsuit\iff b^{\clubsuit}=\heartsuit\]

OpenStudy (anonymous):

x = 4 ans

OpenStudy (anonymous):

and so step one is to write \[\log_3(x-4)=-3\iff x-4=3^{-3}=\frac{1}{3^3}=\frac{1}{27}\] algebra from there on in